# Forecasting and Manipulating the Forecasts of Others

### 18 Mar 2026 v2

@@ Abstract
In strategic environments with private information, evaluating a change in policy requires predicting how the equilibrium responds—but when actions reshape opponents’ signals, each agent’s optimal response depends on an infinite hierarchy of beliefs about beliefs [28] that has resisted exact analysis for four decades. We provide the first exact equilibrium characterization of finite-player continuous-time linear quadratic games with endogenous signals. Conditioning on primitive Brownian shocks rather than the physical state—a dynamic analogue of Harsanyi’s common-prior construction—collapses the belief hierarchy onto deterministic impulse response maps, reducing Nash equilibrium to a deterministic fixed point with no truncation and no large-population limit. The characterization yields an explicit information wedge that prices the marginal value of shifting opponents’ posteriors. The wedge vanishes precisely when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters, and provides a closed-form mapping from information primitives to equilibrium outcomes.
@@ 1 Introduction
In strategic environments with private information, evaluating a change in policy—a disclosure mandate, a shift in the precision of public communication, a redesign of market transparency—requires understanding how the equilibrium responds, not merely how agents behave under existing rules. When a regulator alters the information structure, each agent’s optimal action depends on their beliefs about the state and on their beliefs about how others are filtering, and both change with the policy. The filtering environment and the strategic equilibrium are jointly determined.
This joint determination is the central obstacle. Consider a central bank shifting from quarterly to real-time disclosure of its output gap estimates. Each firm’s pricing decision depends on its belief about the gap and on its belief about how other firms filter the bank’s announcements. Under the new policy, signal precision changes; firms filter differently, revise their forecasts of one another’s pricing, and equilibrium inflation dynamics shift in tandem. Existing tools cannot trace this chain because the Townsend hierarchy makes the mapping from signal precision to equilibrium outcomes intractable. The characterization in this paper makes it explicit.
The same obstruction arises across a wide range of fields. In market microstructure, prices aggregate dispersed information and trading itself is informative. Dynamic information design faces the same difficulty: the mapping from signal structures to equilibrium outcomes becomes intractable once filtering and equilibrium interact. Contract theory encounters it when multiple agents exert unobservable effort on a shared outcome and draw inferences from it. So does decentralized control of autonomous vehicle fleets and multi-robot systems with misaligned operators, wherever agents observe a shared state through private channels and actions reshape what others sense. Existing approaches to the decentralized control problem typically assume cooperative teams or a common information base; the framework here requires neither.
Townsend [28] crystallized the difficulty: in a game with private signals, each agent must forecast others’ forecasts of the state, generating an infinite hierarchy of conditional expectations. Sargent [26] showed that even in linear–Gaussian environments the hierarchy prevents finite-dimensional recursive analysis, because each agent’s action enters the state dynamics, entangling every agent’s past actions, beliefs, and signals in a way the Kalman filter cannot close.
@@
The conceptual move parallels Harsanyi’s [10] resolution of the belief hierarchy in incomplete-information games. Harsanyi replaced an infinite regress of beliefs about beliefs with a common prior over a type space from which all higher-order beliefs follow by conditioning. We replace the Townsend regress with the common probability space of primitive shocks $W=(W^{0},\ldots,W^{n})$, from which all higher-order conditional expectations follow by linear projection onto deterministic impulse response maps with no truncation. The analogy is imperfect in a revealing way: Harsanyi’s types are drawn once and the information partition is fixed, whereas here noise unfolds continuously and actions reshape signals, so the closure requires not just a common prior but a fixed point—the maps governing how projections evolve must be consistent with the strategy profile they induce (Figure 1).
In the single-agent problem the conditional mean of the state is a sufficient statistic, but the state estimate is a linear functional of the noise estimate, not the reverse. In the multi-agent setting the state estimate fails to close the equilibrium because it entangles exogenous shocks with endogenous responses—precisely the Townsend hierarchy. Estimating the noise path avoids this entanglement. We call the path $\widehat{W}_{t}^{i}(\cdot):=\mathbb{E}[W_{\cdot}\mid\mathcal{F}_{t}^{i}]$ the player’s noise-state; optimal controls are affine functionals of it, just as single-agent controls are affine in the state estimate.
The main results establish that the best response amongst all admissible actions to any profile of linear strategies is itself linear. The equilibrium characterization produces an explicit information wedge $\mathcal{V}^{i}_{t}$, which prices the marginal value to player $i$ of shifting opponents’ posteriors. It vanishes identically when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters. The propagation $\Delta P^{i}\to\Delta\widetilde{X}^{i}\to\Delta\mathcal{V}^{i}_{t}\to\Delta D^{i}$ provides an explicit mapping from information primitives to equilibrium outcomes that a model with fixed filtering cannot deliver: because information enters through precision paths and affects the equilibrium only through deterministic impulse response maps, a principal can trace how disclosure precision propagates into equilibrium payoffs without simulating higher-order beliefs.
@@ 1.1 Related Literature
The joint determination of filtering and equilibrium has been approached from several directions, each achieving tractability by restricting or removing the informational externality that makes the problem hard.
@@
The infinite regress of “forecasting the forecasts of others” was identified by Townsend [28] and shown by Sargent [26] to resist finite-dimensional recursive analysis even in linear–Gaussian environments. Morris and Shin [20] demonstrated that the social value of public information can be negative when agents overweight common signals—a static manifestation of the belief-manipulation channel that arises dynamically through the information wedge $\mathcal{V}^{i}_{t}$. Our Proposition 3.6 provides a dynamic counterpart: Section 4 shows that nearly all welfare gains from pooling come from eliminating the strategic channel rather than improving state estimation, consistent with the Morris–Shin intuition that the coordination externality, not statistical efficiency, drives welfare losses.
@@
A substantial literature solves dispersed-information models in the frequency domain [9, 14, 24], culminating in Huo and Takayama [12], who obtain finite-state representations for equilibrium aggregates under exogenous ARMA signals. These methods succeed because the filtering environment is invariant to changes in control laws—the spectral factorization can be solved independently of the policy. When controls enter the state dynamics the transfer function depends on the policy, breaking this separation; the present paper addresses this complementary case.
@@
A powerful technique in decentralized control conditions on a common information base shared by all agents to obtain a recursive reduction [21]. The approach handles asymmetric information but resolves it through a coordinator’s beliefs; the channel by which player $i$’s action shifts player $k$’s posterior, which shifts $k$’s action, which shifts $i$’s signal, is not captured.
@@
The decentralized decision problem with shared objectives originates with Marschak and Radner [19]; Radner [23] established certainty equivalence and linear optimal policies for linear quadratic gaussian (LQG) teams. Team-theoretic solutions bypass the belief hierarchy because a social planner with aligned objectives solves a single-agent control problem—the fixed-point structure of Section 3 never arises. The information wedge $\mathcal{V}^{i}_{t}$ is an equilibrium object: it prices the value of shifting an opponent’s posterior, which requires distinct objectives to be well-defined.
@@
Mean-field games take large-population limits in which no single agent’s action measurably affects any other agent’s signal [6, 11]. Carmona and Delarue [7] provide a comprehensive two-volume treatment including LQG specifications with common noise. The bilateral informational externality is absent by construction in the mean-field limit; $\mathcal{V}^{i}_{t}$ is a finite-$n$ object that vanishes in any limit where individual actions become negligible.
@@
Kyle [15], Back [2], and Foster and Viswanathan [8] solve strategic trading problems with private information, each exploiting special structure—a single informed trader, correlated-signal symmetry—to close the belief hierarchy through model-specific conjectures. Sannikov [25] solves a continuous-time principal–agent problem by showing that the agent’s continuation value serves as a sufficient statistic for the principal’s problem, yielding a recursive characterization despite the agent’s private action. The technique relies on the bilateral structure: the principal’s belief about a single agent’s hidden action can be summarized by a one-dimensional state variable, and the agent has no beliefs about other agents to forecast. With multiple agents exerting unobservable effort on a shared outcome—the moral-hazard analogue of our setting—each agent draws inferences from the shared output about others’ effort, generating the same belief hierarchy that arises here. The noise-state framework extends naturally to this setting: each agent’s hidden effort enters the output dynamics exactly as controls enter the state in (2.1), and the information wedge prices the value of shifting the principal’s and co-agents’ posteriors about one’s effort. The present paper provides a general framework that nests these as special cases: in the Kyle–Back embedding (Section 5), the information wedge corresponds to the strategic component of price impact, and the characterization extends to multi-trader settings where guess-and-verify breaks down.
@@
Kamenica and Gentzkow [13] established Bayesian persuasion; Bergemann and Morris [3, 4] introduced Bayes correlated equilibrium and survey the field, identifying dynamic decentralized settings as a central open problem. Bergemann, Heumann, and Morris [5] provide precision comparative statics in static coordination games analogous to the propagation $\Delta P^{i}\to\Delta\mathcal{V}^{i}_{t}\to\Delta J^{i}$ derived here. The difficulty in extending these results dynamically is that filtering and equilibrium interact; our characterization provides the equilibrium mapping that dynamic information design requires as an input.
Rational inattention.
Sims [27] introduced rational inattention; Maćkowiak and Wiederholt [17] showed that firms optimally attend to idiosyncratic over aggregate conditions, generating sluggish price adjustment (see Maćkowiak, Mat́ejka, and Wiederholt [18] for a survey). These models take the strategic environment as given when solving the attention problem; Corollary 3.7 and the FOC (C.9) extend the framework to settings where the marginal value of precision depends on the equilibrium through the information wedge.

### 23 Mar 2026 v3

@@ Abstract
When actions reshape opponents’ signals, each agent’s optimal response depends on an infinite hierarchy of beliefs about beliefs (Townsend, 1983) that has resisted exact analysis for four decades. We provide the first exact equilibrium characterization of finite-player continuous-time linear–quadratic games with endogenous signals. Estimating primitive shocks rather than the state collapses the belief hierarchy onto deterministic impulse-response maps, reducing Nash equilibrium to a deterministic fixed point with no truncation and no large-population limit. The characterization yields an information wedge pricing the marginal value of shifting opponents’ posteriors; it vanishes precisely when signals are exogenous. In a two-player benchmark, nearly all welfare gains from pooling information come from eliminating bilateral belief manipulation rather than from improving state estimation. This strategic channel also governs optimal precision allocation: a planner should concentrate information on the efficient player, since arming the inefficient player maximizes the arms race while starving them of information collapses it.
@@ 1 Introduction
Dynamic games with dispersed private information pervade economics: firms infer competitors’ pricing from market outcomes, traders learn from the order flow their own trades generate, central banks know their announcements reshape private agents’ signals. In each setting, actions feed back into the information environment, generating an infinite hierarchy of beliefs that Townsend [32] identified and Sargent [29] showed resists finite-dimensional analysis even in linear–Gaussian environments. Bergemann and Morris [6] identify dynamic decentralized information design as a central open problem. Four decades after Townsend, this paper resolves the hierarchy: estimating primitive shocks rather than the state collapses the infinite regress onto deterministic impulse-response maps, reducing Nash equilibrium to a deterministic fixed point. No truncation or large-population limit is needed: a researcher who specifies payoffs, a state equation, and a signal structure gets back equilibrium impulse-response functions; the belief hierarchy is handled internally.
@@
The conceptual move parallels Harsanyi’s [12] resolution of the belief hierarchy in incomplete-information games: just as a common prior over a type space closes the static regress, the common probability space of primitive shocks closes the dynamic one, with all higher-order conditional expectations following by linear projection onto deterministic impulse-response maps. The difference is that Harsanyi’s types are drawn once while noise unfolds continuously and actions reshape signals, so the closure requires a fixed point: the maps governing belief updating must be consistent with the strategy profile they induce (Section 4).
The resolution produces a new object—the information wedge $\mathcal{V}^{i}_{t}$—that prices the marginal value of shifting opponents’ posteriors. If player $i$ increases effort by a small amount, the state drifts, opponent $k$’s signal changes, $k$ revises their posterior and adjusts their action, and the wedge is the present value of that entire chain. It vanishes precisely when signals do not depend on actions, delineating the boundary where strategic belief manipulation matters. The wedge organizes four findings, ordered by their economic content.
Welfare decomposition.
In a two-player game with opposing targets, the simplest setting exhibiting the full bilateral externality, nearly all welfare gains from pooling information come from eliminating the channel by which agents manipulate each other’s posteriors, not from improving state estimation (Section 2). The estimation benefit of pooling is identical whether agents cooperate or compete; the order-of-magnitude difference isolates the strategic externality.
Information starvation.
Allowing asymmetric effort costs, optimal precision allocation under competition is sharply asymmetric and in the opposite direction from what strategic intuition suggests. One might expect a planner to balance precision to limit manipulation—concentrating information arms the efficient player, so strategic restraint should push toward equality. Instead, concentrating information on the efficient player collapses the arms race, while shifting it toward the inefficient player maximizes bilateral waste (Section 2.5).
Separation failure.
Certainty equivalence fails: with multiple agents and endogenous signals, improving one player’s information changes their policy rule itself, not merely their state estimate, because the optimal action depends on what opponents have and have not learned. In any policy evaluation exercise—central bank transparency, disclosure regulation, market structure reform—the information structure cannot be treated as a parameter that shifts beliefs while leaving decision rules fixed; the wedge couples the two, and the equilibrium must be recomputed. The obstruction arises wherever actions enter the state and signals are not fully revealing—the regime Grossman and Stiglitz [10] showed is generic when information acquisition is costly.
Exogenous-signal reduction.
Any model that treats signals as exogenous, including mean-field limits, frequency-domain solutions with fixed spectral densities, and team-theoretic benchmarks, is implicitly assuming away the strategic channel that drives the first three findings. The wedge makes this precise: when controls do not enter the state dynamics, $\mathcal{V}^{i}_{t}\equiv 0$ for all $i$ and $t$, and the equilibrium reduces to $n$ independent single-agent LQG problems against a common exogenous state (Corollary 4.3).
Section 2 previews the economic consequences in a two-player game; Section 4 develops the general theory.
@@ 1.1 Related Literature
The joint determination of filtering and equilibrium has been approached from several directions; each achieves tractability by restricting or removing the informational externality that the present paper resolves exactly.
@@
The infinite regress of “forecasting the forecasts of others” was identified by Townsend [32] and shown by Sargent [29] to resist finite-dimensional recursive analysis even in linear–Gaussian environments. Morris and Shin [23] demonstrated that the social value of public information can be negative when agents overweight common signals; Angeletos and Pavan [1] characterized the efficient use of information in static coordination games, showing that welfare losses depend on the gap between equilibrium and socially optimal degrees of coordination. Woodford [34] showed that imperfect common knowledge generates persistent real effects of monetary policy through higher-order expectations, with the infinite hierarchy truncated by assuming exogenous private signals. Section 2 provides a dynamic counterpart: the information wedge $\mathcal{V}^{i}_{t}$ prices the strategic channel through which actions reshape beliefs, and nearly all welfare gains from pooling come from eliminating this channel rather than from improving state estimation.
@@
A substantial literature solves dispersed-information models in the frequency domain [11, 16, 27], culminating in Huo and Takayama [14], who obtain finite-state representations for equilibrium aggregates under exogenous ARMA signals. These methods succeed because the filtering environment is invariant to changes in control laws—the spectral factorization can be solved independently of the policy. When controls enter the state dynamics, the signal process depends on the policy, the endogenous-information case that Huo and Takayama [14] show cannot admit a finite-state representation in general. The noise-state technique resolves this case exactly: the belief hierarchy collapses onto a deterministic fixed point with no finite-state approximation required.
@@
The common-information approach conditions on a shared information base to obtain a recursive reduction [24]. The approach handles asymmetric information but resolves it through a coordinator’s beliefs; the channel by which player $i$’s action shifts player $k$’s posterior, which shifts $k$’s action, which shifts $i$’s signal, is not captured.
@@
The decentralized decision problem with shared objectives originates with Marschak and Radner [22]; Radner [26] established certainty equivalence and linear optimal policies for Linear–Quadratic–Gaussian (LQG) teams. Team-theoretic solutions bypass the belief hierarchy because a social planner with aligned objectives solves a single-agent control problem—the fixed-point structure never arises. The information wedge $\mathcal{V}^{i}_{t}$ displayed in Section 2 is an equilibrium object: it prices the value of shifting an opponent’s posterior, which requires distinct objectives to be well-defined.
@@
Mean-field games take large-population limits in which no single agent’s action measurably affects any other agent’s signal [7, 13]. Carmona and Delarue [8] provide a comprehensive two-volume treatment including LQG specifications with common noise. The bilateral informational externality is absent by construction in the mean-field limit; the information wedge $\mathcal{V}^{i}_{t}$ of Section 2 is a finite-$n$ object that vanishes in any limit where individual actions become negligible.
@@
Kyle [18], Back [3], and Foster and Viswanathan [9] solve strategic trading problems with private information, each exploiting special structure (a single informed trader, correlated-signal symmetry) to close the belief hierarchy through model-specific conjectures. Sannikov [28] solves a continuous-time principal–agent problem by showing that the agent’s continuation value serves as a sufficient statistic for the principal’s problem, yielding a recursive characterization despite the agent’s private action. The technique relies on the bilateral structure: the principal observes output directly, so the agent’s deviation changes her payoff but not her posterior, and the wedge $\mathcal{V}^{i}_{t}$ is absent. With multiple agents exerting unobservable effort on a shared outcome, each agent infers others’ effort from the shared output, the noise-state path replaces the scalar continuation value, and the belief hierarchy reappears. The present paper provides a general framework that nests these as special cases: in the Kyle–Back embedding (Section 5), the information wedge corresponds to the strategic component of price impact, and the characterization extends to multi-trader settings where guess-and-verify breaks down.
Rational inattention.
Sims [31] introduced rational inattention; Maćkowiak and Wiederholt [21] showed that firms optimally attend to idiosyncratic over aggregate conditions, generating sluggish price adjustment (see Maćkowiak, Mat́ejka, and Wiederholt [20] for a survey). These models take the strategic environment as given when solving the attention problem; the first-order conditions derived here (Section D.3) extend the framework to settings where the marginal value of precision depends on the equilibrium through the information wedge.
@@
Kamenica and Gentzkow [15] established Bayesian persuasion; Bergemann and Morris [5, 6] introduced Bayes correlated equilibrium and survey the field, identifying dynamic decentralized settings as a central open problem. Vives [33] develops the interplay between information acquisition and competition in financial markets that the Kyle–Back embedding of Section 5 operationalizes dynamically. Bergemann, Heumann, and Morris [4] provide precision comparative statics in static coordination games where the effect of signal precision on payoffs admits direct analysis; dynamically, the information wedge couples the value of precision to the equilibrium, so that filtering and control no longer separate. The equilibrium characterization derived here (Section 4) provides the mapping from signal structures to equilibrium outcomes that dynamic information design requires as an input.
2 Bilateral belief manipulation in a two-player game
Before developing the general framework, we display its consequences in the simplest game exhibiting the full bilateral externality: two symmetric players with scalar state, opposing targets, and equal signal precision. Neither player observes the state directly; each receives a private diffusion signal whose precision $p>0$ is common knowledge. Each player’s control enters the state dynamics and reshapes the opponent’s signal, generating the bilateral informational externality of Townsend [32]. The general theory (Sections 34) collapses this hierarchy onto deterministic impulse-response maps and produces an object, the information wedge $\mathcal{V}^{i}_{t}$, that prices the marginal value of shifting the opponent’s posterior.
2.1 Primitives and objectives
Fix a horizon $[0,T]$. The state $X_{t}\in\mathbb{R}$ satisfies
$dX_{t}=\big(D_{t}^{1}+D_{t}^{2}\big)\,dt+dW_{t}^{0},\qquad X_{0}=x_{0},$ (2.1)
where $W^{0}$ is a standard Brownian motion. Player $i\in\{1,2\}$ does not observe $X$ directly; instead they observe the private diffusion channel
with $(W^{0},W^{1},W^{2})$ mutually independent. Player $i$ chooses $D^{i}$ adapted to their private observation history $\mathcal{F}_{t}^{i}=\sigma(Y_{s}^{i}:s\leq t)$.
Player 1 wants the state near +1+1; player 2 wants it near 1-1. There is no terminal cost in this worked example. The model is invariant under the sign/label symmetry $(X,Y^{1},Y^{2},D^{1},D^{2})\mapsto(-X,Y^{2},Y^{1},-D^{2},-D^{1})$.
[…]
Figure 1: Information-to-action feedback in the worked example. Actions move $X_{t}$ and shift opponents’ beliefs, so equilibrium mean actions depend on precision pp.
2.2 What the general theory delivers
The equilibrium characterization developed in Sections 34 produces three objects that organize the analysis of this game; we describe them informally here and derive them in full generality below.
Noise-state and linear strategies.
All randomness originates in the three primitive Brownian motions $W=(W^{0},W^{1},W^{2})$. Player $i$’s noise-state $\widehat{W}_{t}^{i}(u):=\mathbb{E}[W_{u}\mid\mathcal{F}_{t}^{i}]$ ($0\leq u\leq t$) is the running conditional mean of the entire primitive noise path. In equilibrium, each player’s control takes the form
$D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{i}(u),$ (2.4)
where $\bar{D}^{i}_{t}$ is a deterministic mean action and $D^{i}_{t}(u)$ is a deterministic impulse-response kernel governing how player $i$’s action at time $t$ responds to news about the $u$-th shock. The key structural result (Corollary 4.5) is that the best response to any such strategy profile, over all admissible controls, is itself of this form—so equilibrium reduces to a deterministic fixed point in the maps $(\bar{D}^{i},D^{i})$.
Unresolved impulse response.
The state’s impulse response to the $u$-th shock decomposes into what player $i$ has learned and what remains hidden: $X_{t}(u)=\widehat{X}^{i}_{t}(u)+\widetilde{X}^{i}_{t}(u)$. The unresolved component $\widetilde{X}^{i}_{t}(u)$ is the portion of the $u$-th shock’s effect on the state that player $i$ has not yet resolved—large when the shock is recent or precision low, shrinking as observations accumulate. It serves as the Kalman gain in the noise-state update (Theorem 4.1).
Information wedge.
In the single-agent problem, the backward equation for the costate (the shadow value of the state) is standard. With two agents and endogenous signals, a new term appears: the information wedge
$\bar{\mathcal{V}}^{1}_{t}=P^{2}(t)\int_{0}^{t}\widetilde{X}^{2}_{t}(z)\cdot\bar{H}^{2}_{t}(z)\,dz,$ (2.5)
where $\widetilde{X}^{2}_{t}(z)$ is the unresolved component of the state’s response to the $z$-th shock from player 2’s perspective and $\bar{H}^{2}_{t}(z)$ is the mean belief adjoint—the shadow value to player 1 of having shifted player 2’s estimate of the $z$-th shock (Theorem 4.2). The wedge enters the mean costate equation as $\tfrac{d}{dt}\bar{H}^{X}_{t}=-(\bar{X}_{t}-1)-\bar{\mathcal{V}}^{1}_{t}$, distorting mean actions relative to the single-agent benchmark.
2.3 Computation
Equilibrium is computed by iterating the best-response operator on the deterministic impulse-response system (Section 4), discretized on a uniform grid of $N=40$ points; no Monte Carlo simulation is needed. Parameters: $T=1$, $x_{0}=0$, $r=0.1$, no terminal cost. The Picard residual reaches $10^{-5}$ in $19$ iterations ( ${\approx}\,10$ ms in C++), decaying geometrically. [^1] An interactive browser visualization is available at https://sbabichenko.com/lqg. At longer horizons the binding constraint is numerical resolution (the filtering kernel develops near-singularities) rather than economic.
2.4 Mean actions, separation failure, and the welfare cost of belief manipulation
Since the stochastic integral in (2.4) has zero unconditional mean, $\mathbb{E}[D_{t}^{i}]=\bar{D}^{i}_{t}$ isolates the deterministic tug-of-war over the target $\pm 1$, while the kernel term captures impulse responses to posterior shocks.
Separation failure.
Under perfect information, the mean policy is independent of signal precision. Decentralized information breaks this: precision changes the speed of opponents’ updating, creating an incentive to act on posteriors rather than on the state. The mechanism is the information wedge (2.5): it couples the costate to the unresolved kernel $\widetilde{X}^{2}_{t}(z)$ through the belief adjoint, introducing a dependence of the equilibrium policy on precision that is absent in the single-agent problem. More precisely, in the single-agent case the policy kernel factors as $D^{i}_{t}(u)=S(t)\cdot X_{t}(u)$ for a scalar Riccati gain $S(t)$ independent of signal precision. With two agents, the wedge prevents this factorization: changing $p$ changes the unresolved kernel, which changes the belief adjoint, which changes the policy kernel independently of the state estimate (Proposition 4.6).
The wedge across precision asymmetries.
Figures 34 display the information wedge and its consequences when player 1’s precision is fixed at $p_{1}=3$ and player 2’s precision $p_{2}$ varies.
Information pooling as a Pigouvian intervention.
Figure 5 compares equilibrium costs under private signals with those under a pooled signal of precision $P^{1}+P^{2}$, for both competitive (opposing targets $\pm 1$) and cooperative (common target 0) specifications. Pooling is Pareto-improving in both cases, but the gains differ by an order of magnitude: roughly 0.020.020.050.05 under cooperation versus 0.30.30.50.5 under competition. Since the estimation benefit of pooling enters identically regardless of the target structure, the difference isolates the strategic externality channel priced by $\mathcal{V}^{i}_{t}$: nearly all of the welfare gain from mandatory disclosure comes from eliminating bilateral belief manipulation, not from improving state estimation.
Remark 2.1 (Impulse responses do not depend on targets).
The equilibrium impulse-response maps—how players respond to shocks—do not depend on the targets $\pm 1$. All filtering objects, unresolved components, and policy response maps are identical whether players cooperate or compete; only the deterministic tug-of-war over mean actions sees the strategic structure, through the information wedge.
Figure 5: Equilibrium costs under private signals (dashed) and pooled signals (solid) with $p_{1}=3$ fixed and $p_{2}$ varying. Left: opposing targets ($\pm 1$). Right: common target (0). Pooling is Pareto-improving in both cases, but the gain is an order of magnitude larger under competition, where it eliminates the bilateral belief-manipulation externality priced by $\mathcal{V}^{i}_{t}$. Blue: player 1; red: player 2.
2.5 Optimal precision allocation
Information pooling is a blunt instrument: it merges all signals into a common channel, eliminating some of the bilateral externality but forecloses targeted allocation. Given a fixed precision budget, how should a planner distribute it between the two players?
Setup.
Retain the state dynamics (2.1) and observation channels (2.2) with opposing targets ($\pm 1$), but allow asymmetric precisions $(p_{1},p_{2})$ subject to $p_{1}^{2}+p_{2}^{2}=\bar{P}$ and asymmetric effort costs $(r_{1},r_{2})$ with $r_{1}<r_{2}$, so that player 1 is the more efficient mover of the state. The quadratic budget reflects diminishing returns to concentration. A planner chooses the split to minimize aggregate equilibrium cost; for each candidate allocation we solve the full equilibrium via Picard iteration. Parameters: $\bar{P}=20$, $T=1$, with the fundamental volatility in (2.1) reduced to $\sigma=0.5$.
Figure 6: Equilibrium quantities across the precision split, opposing targets ($\pm 1$), $p_{1}^{2}+p_{2}^{2}=\bar{P}=20$, $\sigma=0.5$, $T=1$. Top: asymmetric effort costs ($r_{1}=0.05$, $r_{2}=0.2$); bottom: symmetric ($r_{1}=r_{2}=0.1$). Left: individual mean control energies; center: total destructive effort (equilibrium solid, full-information dashed); right: mean controls and mean state (equilibrium solid, certainty-equivalent dotted).
What one might expect.
Under cooperation the answer is straightforward: give precision to the efficient player, perhaps sharing some to ease the inefficient player’s burden. Under competition, concentrating precision arms the efficient player to pull the state toward their target, so one might expect the planner to compensate by balancing precision to limit manipulation. The presumed tradeoff is productive efficiency versus strategic restraint.
Information starvation.
The tradeoff is real but the strategic-restraint side is far weaker than expected (Figure 6, top row, $r_{1}=0.05$, $r_{2}=0.2$). Concentrating all precision on the efficient player does shift the mean state further from zero, but the effect on total effort dominates: destructive effort drops by nearly half. The productive-efficiency and strategic-restraint channels point in the same direction: the presumed tradeoff does not exist. The mechanism is information starvation: denying the inefficient player signal access collapses their ability to exert mean control (top left panel: player 2’s effort falls to near zero). With player 2 largely neutralized, player 1 has less reason to manipulate as well, and both players’ effort falls. Conversely, giving all precision to the inefficient player produces the worst outcome: the cheap mover pushes hard against a well-informed opponent who retaliates sharply, maximizing bilateral waste. The planner’s optimal response is not to balance information but to create an information monopoly for the efficient player. The cost of one-sided dominance is a displacement in the mean state; the benefit is the collapse of bilateral waste. The collapse of bilateral waste dominates.
Remark 2.2 (Commitment through incapacity).
The mechanism parallels Schelling’s [30] observation that in a game of chicken, a driver who visibly removes the steering wheel forces the opponent to swerve. Here precision is common knowledge and enters the equilibrium through deterministic kernels, so denying the inefficient player signal access is a credible and observable disarmament: both players benefit not from chosen restraint but from one side’s visible inability to fight.
Separation failure in a picture.
The right panels display separation failure directly: certainty-equivalent mean controls (dotted) are invariant to the precision split, while equilibrium mean controls (solid) slope with the allocation through the belief adjoints. Under symmetric effort costs (bottom row) the pattern is symmetric around the equal split, confirming that the asymmetry above is driven by cost differences interacting with the information channel.